Diffusion Morning Report 10/27/06

Ted Berg tberg at stanford.edu
Fri Oct 27 06:31:28 PDT 2006


Good Morning All,
    Looking OK this morning. LTO is down for broken quartz. Tylan 3 has 
temp issues doing temp cals. Epi has posssible handling issues.
Waiting for AMAT process for new epi. No other issues at this time.

Diffusion Mike,Ray, Ted 	  	  	  	  	 
  	Date: 	10/27/06 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	Up 	  	  	  	 
Tylan 2 	Up 	  	  	  	 
Tylan 3 	down 	temp issues 	  	  	 
Tylan 4 	Up 	  	  	  	 
Tylan 5-BBr3 	Up 	  	  	  	 
Tylan 6-POCl3 	Up 	  	  	  	 
Tylan 7-TEOS2 	Up 	  	8.79/30 	  	 
Tylan 8-SiGe 	Up 	  	6.75/30 	  	 
Tylan 9-Poly 	Up 	  	12.51/30 	  	 
Tylan 10-Nitride 	Up 	  	13.33/30 	  	 
Tylan 11-BPSG-LTO 	down 	broken sheath and injectors 	2.93/30 	  	 
Tylan 12-FGA 	Up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	Up 	  	  	  	 
Thermco 2 	Up 	  	  	  	 
Thermco Poly 	Up 	  	  	  	 
Thermco Nitride 	Up 	  	  	  	 
Epi 	yellow 	possible handling issues 	  	  	 
  	
	
	
	
	
Burn Box 	up 	  	  	  	 
ZYGO 	Up 	  	  	  	 
  	up 	  	   	  	 
GAAS 22 	  	  	  	  	 
GAAS 23 	UP 	  	  	  	 
SiH4 for furnaces 	up 	  	  	  	 

	  	B on line 60 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20061027/488f9c0e/attachment.html>


More information about the diffusion mailing list